OpenStudy (anonymous):

how to solve 2x-y=2 & y=0

OpenStudy (anonymous):

2x-y=2 -y=2-2x y=-2+2x

OpenStudy (anonymous):

or 2x-0=2 2x=2 x=1 ?

OpenStudy (anonymous):

Thank you!